OpenStudy (gabylovesyou):

@satellite73 how do i graph f(x) = the quantity of 9 x squared plus 9 x minus 18, all over 3 x plus 6?

OpenStudy (gabylovesyou):

\[f(x) = \frac{ 9x^2 + 9x - 18 }{ 3x + 6 }\]

OpenStudy (anonymous):

\[f(x)=\frac{9x^2+9x+18}{3x+6}\]

OpenStudy (anonymous):

oh minus ok

OpenStudy (anonymous):

gonna factor and cancel like before

OpenStudy (anonymous):

you got that?

OpenStudy (gabylovesyou):

yes..

OpenStudy (gabylovesyou):

9(x+2)(x-1)

OpenStudy (anonymous):

yeah cancel with the denominator and get \(y=3x-3\)

OpenStudy (gabylovesyou):

ok and then

OpenStudy (anonymous):

since the denominator factors as \(3(x+2)\)

OpenStudy (anonymous):

graph of \(y=3x-3\) a line with slope 3, y intercept -3

OpenStudy (gabylovesyou):

so its or C or D.. what is circled..

OpenStudy (anonymous):

open circle: set \[3x+6=0\] get \[x=-2\]

OpenStudy (anonymous):

plug in \(-2\) to \(y=3x-3\) get \(-9 \)

OpenStudy (gabylovesyou):

so C ?

OpenStudy (anonymous):

yup

OpenStudy (gabylovesyou):

thank you ;)

OpenStudy (anonymous):

yw